OpenStudy (anonymous):

Select the ordered pair that is the solution of this equation: -7x – 4y = 6

OpenStudy (anonymous):

choices are : 2,-2 -2,2 2,2 2,0

OpenStudy (anonymous):

satellite can you please help me on my questions :(

OpenStudy (anonymous):

replace x and y by the choices given, ans see which works. for example the first one give \[-7\times 2-4\times(- 2)=6?\] \[-14+8=6?\] \[-6=6?\]nope

OpenStudy (anonymous):

i just did it on paper and i cant figure it out , its like none of them is right .